Week of 160822

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes




  • local: Andrea S. (SCOD), Michal (ATLAS), Iain (Grid and Compute services), Hervé (Storage), Nilo (Databases), Andrea M. (FTS, MW), Alberto (Monitoring), Vincent (Security)
  • remote: Christoph (CMS), Di (TRIUMF), Dave (FNAL), Eric L. (BNL), Eric Cogneras, Tiju (RAL), Vincenzo (EGI), Dmitri (KIT), Pepe (PIC), Kyle (OSG)

Experiments round table:

  • CMS reports (raw view) -
    • There is small chance that I cannot make it to the meeting (sorry)
    • Rather quiet operations
    • Outage of network link between Wigner und Meyrin did not lead to (obvious) problems
    • Executed a scheduled cleaning of DDM disk (volume several PBs)
    • Short CASTOR intervention during MD of this week is o.k. for CMS (both proposed dates would be fine)

  • ALICE -
    • NTR

  • LHCb reports (raw view) -
    • Activity
      • Monte Carlo simulation, data reconstruction and user jobs on the Grid
    • Site Issues
      • T0:
        • NTR
      • T1:
        • jobs which failed to upload data to GRIDKA-BUFFER (GGUS:123522) and to RAL-BUFFER (GGUS:123521) (In progress)
        • Failing transfers at INFN with 530-an unknown (GGUS:123538) (In progress)

Sites / Services round table:

  • ASGC:
  • BNL: ntr
  • CNAF:
  • EGI: ntr
  • FNAL: ntr
  • GridPP:
  • IN2P3: Possibly nobody from IN2P3-CC can connect today. There is an announcement though: an all-day maintenance downtime is foreseen for September 20th. The planning is not yet finalized. More details will be available one week before the outage.
  • KISTI:
  • KIT: ntr
  • NDGF: Downtime on Thursday 10-14 CEST. Slurm, switch firmware and Dcache upgrade at UiB. Dcache upgrade at several other sub-sites will be made using the same downtime window. Some data unavailable, but should only be for short periods during the window as pool nodes are rebooted. (Possibly wont make it back from another meeting in time to connect. Christian)
  • NL-T1:
  • NRC-KI: could not attend the meeting as everybody is busy getting the site up after the downtime. ETA is tomorrow, checking all bits and updating software/firmware today (Eygene)
  • OSG: ntr
  • PIC: ntr
  • RAL: ntr
  • TRIUMF: ntr

  • CERN computing services: ntr
  • CERN storage services:
    • FTS:
      • 2 VMs down last week because of hypervisors issues: 1 on Tue reported by LHCb (GGUS:123413) and the other one on Thu (OTG:0032450). In both cases the transfers running on the VMs were aborted.
  • CERN databases:
    • CMSONR suffered some connectivity problems during the week-end because a couple of CMS DNS servers were not pingable any more. CMS provided an alternative list that we added to the the database servers OS configuration.


Edit | Attach | Watch | Print version | History: r17 < r16 < r15 < r14 < r13 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r17 - 2016-08-22 - AndreaSciaba
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback